# Modeling AdaGrad, RMSProp, and Adam with Integro-Differential Equations

### v1

@@ 5 Conclusion
Our continuous models establish a direct connection between the discrete implementations of these algorithms and their continuous counterparts. The nonlocal integral terms naturally encapsulate the algorithms’ memory and adaptive behaviors, allowing us to represent the influence of past gradients on current updates explicitly. Through extensive numerical simulations, we have demonstrated that these continuous formulations accurately replicate the convergence dynamics and behaviors of the original algorithms across various settings, including different learning rates and parameter configurations. Furthermore, as the learning rate is reduced, the similarity between the continuous and discrete models increases, further confirming the effectiveness of the continuous representation in capturing the behavior of the discrete algorithms.
@@ Acknowledgments
Appendix
Appendix A Proofs
In this section, we present detailed proofs for Propositions 1 to 3.
Proof of Proposition 1:
Proof.
By applying Assumption (2), the evolution of the accumulated gradients can be described by the differential equation:
whose solution (for $t>0$) is:
Using Assumption (2) again, we derive the integro-differential equation that governs the updated parameters:
Proof of Proposition 2:
Proof.
By applying Assumption (2), the evolution of the squared gradient moving average is governed by the differential equation:
$\dot{v}+\frac{1-\beta_{2}}{\alpha}v=\frac{1-\beta_{2}}{\alpha}\partial^{i}f(\theta)\partial_{i}f(\theta)\,.$ (7)
Rewriting this equation yields:
$\frac{{\rm d}}{{\rm d}t}\left[e^{\frac{1-\beta_{2}}{\alpha}t}v\right]=\frac{1-\beta_{2}}{\alpha}e^{\frac{1-\beta_{2}}{\alpha}t}\partial^{i}f(\theta)\partial_{i}f(\theta)\,,$ (8)
which leads to the solution (for $t>0$):
Finally, by applying Assumption (2), the parameter updates follow the integro-differential equation:
Proof of Proposition 3:
Proof.
Applying Assumptions (2) and (1) and performing some algebraic manipulations yields:
where
Appendix B Additional Numerical Simulations
In this appendix, we present additional simulations for the nonlocal continuous models and compare them with their conventional counterparts. Unlike the simulations presented in Section 4.2, this section focuses on the convergence behavior of the Mean Squared Error (MSE) loss function for the function $y(x)=2x$, where the target parameter value to be obtained is 2.
The settings for these simulations were as follows:
  • Figure 15: Simulations were performed with RMSProp ($\beta=0.9$) and Adam ($\beta_{1}=0.9,\beta_{2}=0.99$) optimizers, using learning rates of 0.1 and 0.01, respectively. The parameter $\theta_{0}$ was initialized to 0. To ensure computational efficiency, simulations were limited to 60 k-iterations for a learning rate of 0.1 and 300 k-iterations for a learning rate of 0.01, which was sufficient to capture the desired behavior.
  • Figure 16: Simulations were conducted with RMSProp ($\beta=0.99$) and Adam ($\beta_{1}=0.99,\beta_{2}=0.999$) optimizers, again using learning rates of 0.1 and 0.01, respectively. The parameter $\theta_{0}$ was initialized to 0. To maintain computational efficiency, simulations were capped at 80 k-iterations for a learning rate of 0.1 and 500 k-iterations for a learning rate of 0.01.
Figures 15 and 16 demonstrate that the numerical simulations of the integro-differential equations closely match the results of the discrete models, regardless of the optimizer used, reaffirming the validity of the continuous models. As expected, the AdaGrad model is more computationally intensive to reach a loss of 0, while RMSProp achieves convergence more quickly. An interesting observation is that, for a learning rate of 0.1, the nonlocal models accurately reproduce the bump around 35 k-iterations, which gradually diminishes as the learning rate decreases. Moreover, it is evident that as the learning rate is lowered, the continuous model more closely approximates the behavior of the discrete model.
Figure 15: MSE Loss Convergence for Different Solvers. The plots show the convergence behavior of the Mean Squared Error (MSE) loss for various optimization algorithms (AdaGrad, Adam, and RMSProp) along with their nonlocal continuous versions, evaluated at two different learning rates (0.1 and 0.01). The left plot shows results for a learning rate of 0.1, while the right plot presents those for a learning rate of 0.01. The configurations used are RMSProp with $\beta=0.9$ and Adam with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. In general, lower learning rates tend to produce smoother convergence curves, whereas higher learning rates can accelerate initial convergence but often lead to more fluctuations.
Figure 16: MSE Loss Convergence for Different Configuration Solvers. The plots display the convergence of the Mean Squared Error (MSE) loss for various optimization algorithms (AdaGrad, Adam, and RMSProp) and their nonlocal continuous counterparts, using two different learning rates (0.1 and 0.01). The configurations include RMSProp with $\beta=0.99$ and Adam with $\beta_{1}=0.99$ and $\beta_{2}=0.999$. The left plot shows results for a learning rate of 0.1, while the right plot corresponds to a learning rate of 0.01. Lower learning rates typically produce smoother convergence curves, whereas higher learning rates may lead to faster initial convergence but can also cause more pronounced fluctuations.

### v2

@@ 6 Conclusion
Our continuous models establish a direct connection between the discrete implementations of these algorithms and their continuous counterparts. The nonlocal integral terms naturally encapsulate the algorithms’ memory and adaptive behaviors, allowing the explicit representation of the influence of past gradients on current updates, Propositions 1 to 3.
Moreover, we study the convergence and stability properties of these continuous models. Theorem 1 establishes exponential convergence while revealing that memory effects slow down the rate. Under $L$-smoothness and strong convexity, we construct a Lyapunov function that ensures both convergence and stability of the equilibrium point $\theta^{\ast}$, yielding $\|\theta(t)-\theta^{\ast}\|\leq\|\theta(0)-\theta^{\ast}\|e^{-m\mu_{0}t}\,$, with an explicit decay constant $\mu_{0}$ that decreases with both the memory magnitude $\|\nu\|$ and the initial optimization gap. The distinction between AdaGrad and RMSProp arises naturally through the parameter $\nu$.
To separate accumulation from forgetting, Theorem 2 introduces the functional $E_{\rho}=\Phi+\rho\,M_{\nu}[g^{2}]$ and proves a closed dissipation inequality $\dot{E}_{\rho}\leq-\kappa_{0}E_{\rho}$, where $\kappa_{0}=\min\{m\mu_{0},\lambda\xi\}$. The memory obeys $\dot{M}_{\nu}[g^{2}]=-\xi\,M_{\nu}[g^{2}]+g^{2}$, so $\xi M_{\nu}[g^{2}]$ acts as forgetting and $g^{2}$ acts as accumulation. Hence RMSProp ($\xi=\lambda$) yields exponential decay of $E_{\rho}$, both the memory and the objective gap via $\kappa_{0}$, whereas AdaGrad ($\xi=0$) is a pure accumulator, $\dot{E}=0$, ensuring monotone descent —via Theorem 1— but no exponential rate. Regarding the rate, $m\mu_{0}$ decreases as the memory magnitude $\|\nu\|$ grows (more memory implies smaller $\mu_{0}$), while $\lambda\xi$ increases with $\lambda$. Consequently, $\kappa_{0}$ increases with weaker memory (smaller $\|\nu\|$) up to the ceiling $m\mu_{0}$, but decreases with stronger memory (larger $\|\nu\|$). In particular, for AdaGrad $\kappa_{0}=\min\{m\mu_{0},0\}=0$, hence no exponential rate.
For Adam, Theorem 3 establishes both convergence and stability through the Lyapunov functional $V=\Phi+\frac{\alpha}{2\lambda_{1}}\,\|M_{\nu_{1}}[g]\|^{2}.$ Stability follows from the Lyapunov framework itself, while convergence arises under the assumptions $h(t)=\frac{{\rm d}}{{\rm d}t}\log\alpha(t)\leq C_{\alpha}<2\lambda_{1}$ and $\alpha(t)\geq\alpha^{\ast}>0$. By applying LaSalle’s invariance principle to $V$, we obtain global convergence, namely $\theta(t)\to\theta^{\ast}$ and $\|M_{\nu_{1}}[g](t)\|\to 0$. Once convergence is guaranteed, using an inequality of the damped oscillator of $V$, we obtain exponential convergence to the minimizer $\theta^{*}$.
The nonconvex analysis mirrors this picture. Lemma 11 ensures that, for AdaGrad and RMSProp, we have $\theta(t)\to\theta^{\ast}$ as $t\to\infty$ by an application of Barbălat’s lemma, since $\nabla f(t)\to 0$ in the limit. Building on Lemma 11, Theorem 4 guarantees exponential decay under the Polyak–Łojasiewicz (PL) condition, and finite, polynomial or exponential convergence under the more general Kurdyka–Łojasiewicz (KL) framework. Specifically, if the KL exponent satisfies $\theta\in(0,1/2)$, convergence occurs in finite time, whereas for $\sigma=1/2$ it becomes exponential. Finally, when $\sigma\in(1/2,1)$, the rate transitions to a polynomial regime, $O(t^{-\tfrac{1}{1-2\sigma}}),$ all of which hold locally, in a neighborhood of the stationary point. Following the same reasoning and using the same functionals as in Theorems 2 and 3, Theorems 5 and 6 extend the energy functionals $E_{\rho}$ (for AdaGrad/RMSProp) and $V$ (for Adam) to the nonconvex setting. These results ensure that $\theta(t)$ stabilizes toward the critical set —via LaSalle’s invariance principle in Theorem 6—, and that exponential convergence holds under the PL condition, while KL-type local rates govern the general nonconvex case. In this regime, both the gradient norms and Adam’s momentum terms vanish asymptotically within the attraction basin of a local minimizer $\theta^{\ast}$, assuming such a minimizer exists. However, these results do not assert the uniqueness of the limit point, as multiple local minima may coexist.
Through extensive numerical simulations, we have demonstrated that these continuous formulations accurately reproduce the convergence dynamics and behaviors of the original algorithms across a wide range of settings, including different learning rates and parameter configurations. Furthermore, as the learning rate decreases, the similarity between the continuous and discrete models increases, confirming the effectiveness of the continuous representation in capturing the behavior of the discrete algorithms. The simulations also align closely with the theoretical predictions: in AdaGrad, trajectories approach minimizers in the convex case (critical points in the nonconvex case) while the accumulated memory saturates and the effective step shrinks, flattening the curves; in RMSProp, the memory decay exponentially, with slower rates as the memory becomes stronger ($\beta\to 1$); and in Adam, both moments damp and the parameters stabilize, with longer transients under stronger memory.
@@ Acknowledgments
Author Contributions
Carlos Heredia was solely responsible for the conception, development, implementation, and writing of this manuscript.
Funding
The author declares that no funding was received for this work.
Appendix A Nonconvex Simulation. Case $\theta^{*}=0$.
This section illustrates the simulations for the $\theta^{*}=0$ case of the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$. Note that this case is trivial: since the initial condition is $\theta_{0}=0$, the system starts at the critical (maximum) point and therefore remains there, as shown in Fig. 15. As can be observed in the non-convex plots, any small perturbation causes the system to converge to one of the local minima.
On the other hand, since all the plots for AdaGrad, RMSProp, and Adam are identical, with all trajectories remaining at the minimum $\theta_{0}=0$, we only display the first one: AdaGrad. The others can be found in the simulations/figures folder of the GitHub repository.
Appendix B Simulation Figures
In this section, we present the simulation figures for both the convex and nonconvex cases.
B.1 Convex Cases
Figure 1: Convergence of $\theta(t)$ using the first-order nonlocal continuous AdaGrad method. The plot illustrates the convergence trajectories for minimizing the function $(\theta-4)^{2}$ using two different learning rates: $0.1$ (left) and $0.01$ (right). At a higher learning rate ($0.1$), the algorithm rapidly converges to the target value $\theta=4$, stabilizing in under 1,500 k-iterations. With a lower learning rate ($0.01$), the convergence is more gradual, reaching the target around 100,000 k-iterations.
Figure 3: Convergence of $\theta(t)$ using the first-order nonlocal continuous RMSProp method. The plot shows the convergence to the minimum value of $\theta=4$ for the convex function $(\theta-4)^{2}$. The left subplot corresponds to a learning rate of 0.1, while the right subplot uses a learning rate of 0.01. With a learning rate of 0.1, $\beta=0.0$ exhibits more noticeable oscillations, whereas $\beta=0.9$ and $0.99$ converge more smoothly. For a learning rate of 0.01, a slight destabilization occurs for $\beta=0.9$ as the solution approaches the final result, caused by the numerical method. At higher learning rates, slight differences can be observed between the models: for $\beta=0.0$, the oscillations in the discrete case start immediately upon reaching the minimum, whereas in the continuous case, they take a few k-iterations to begin. On the other hand, for $\beta=0.99$, the curve is less pronounced in the discrete model compared to the continuous one.
Figure 4: Convergence trajectories of $v(t)$ for the first-order nonlocal continuous RMSProp method. This plot illustrates the convergence of the squared gradient moving average $v(t)$. The left subplot represents a learning rate of 0.1, while the right subplot corresponds to a learning rate of 0.01. For $\beta$ values of 0.0 and 0.9, a slight initial bump is noticeable, but both eventually decay towards zero. The main differences between the continuous and discrete models appear at a learning rate of $0.1$, where the values are slightly higher for $\beta=0.9$ and $\beta=0.99$, and for $\beta=0.0$ a small bump is observed instead of a direct descent.
Figure 5: Convergence of $\theta(t)$ using the first-order nonlocal continuous Adam method. The plot illustrates the convergence trajectories of $\theta$-values for the first-order nonlocal continuous Adam model under different parameter settings of $\beta_{1}$ and $\beta_{2}$ with two distinct learning rates ($0.1$ and $0.01$). For $\beta_{1}=0.9$ and a learning rate of 0.1, a noticeable oscillation is observed, requiring a longer time to stabilize at the minimum value. This behavior improves when the learning rate is reduced to 0.01.
Figure 6: Convergence trajectories of $m(t)$ for the first-order nonlocal continuous Adam optimizer. The plot shows the evolution of the first moment $m(t)$ over time, scaled by the learning rate. With a higher learning rate ($\alpha=0.1$), the values of $m(t)$ exhibit certain oscillations, whereas a lower learning rate ($\alpha=0.01$) results in a smoother and slower convergence.
Figure 7: Convergence trajectories of $v(t)$ for the first-order nonlocal continuous Adam optimizer. The plot illustrates the convergence of the squared gradient moving average $v(t)$. At a higher learning rate ($\alpha=0.1$), the peaks of $v(t)$ vary according to the parameter settings, with $\beta_{1}=0.0$ and $\beta_{2}=0.99$ producing the most significant peak and the slowest rate of decay. When the learning rate is reduced to $0.01$, all curves display a more pronounced initial peak followed by a gradual decline.
B.2 Nonconvex Cases
(a) Convergence of $\theta(t)$ with $\theta_{0}=0.1$ for AdaGrad. The nonlocal method and discrete AdaGrad converge monotonically towards $\theta^{*}\approx 1$. With $\alpha=0.1$, the nonlocal approximation is slightly smoother; with $\alpha=0.01$, both trajectories overlap almost completely, although they require many more iterations.
(b) Convergence of $\theta(t)$ with $\theta_{0}=-0.1$ for AdaGrad. Symmetric behavior with respect to the positive case: monotonic convergence to $\theta^{*}\approx-1$. At $\alpha=0.1$, the nonlocal model is slightly slower at the beginning; at $\alpha=0.01$, both curves essentially coincide throughout the entire interval.
Figure 8: AdaGrad numerical simulations. These two plots show the evolution of the trajectory of $\theta(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ with two initial conditions $\theta_{0}=\pm 0.1$ for AdaGrad.
(a) Gradient trajectories for nonlocal and discrete AdaGrad dynamics with initial value $\theta_{0}=0.1$. Gradient trajectories with $\theta_{0}=0.1$. For $\alpha=0.1$, the nonlocal model reaches a slightly higher plateau and shows a smoother transition than discrete AdaGrad; for $\alpha=0.01$, both trajectories overlap almost completely throughout $t/\alpha$.
(b) Gradient trajectories for nonlocal and discrete AdaGrad dynamics with initial value $\theta_{0}=-0.1$. Gradient trajectories with $\theta_{0}=-0.1$. Results analogous to the case $\theta_{0}=0.1$: at $\alpha=0.1$ the nonlocal term is smoother and reaches a slightly higher plateau; at $\alpha=0.01$ the agreement with the discrete scheme is practically exact.
Figure 9: Gradient trajectories of AdaGrad. These two plots show the evolution of the trajectory of the gradient $G(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ with two initial conditions $\theta_{0}=\pm 0.1$ for AdaGrad.
(a) Convergence trajectories of $\theta(t)$ with $\theta_{0}=0.1$ for the nonlocal and discrete models. At $\alpha=0.1$, high momentum ($\beta=0.99$) leads to pronounced overshoot/oscillations; $\beta=0$ and $0.9$ are smoother. At $\alpha=0.01$, all settings converge smoothly with ordering of speeds $\beta=0.99>0.9>0$. Nonlocal and discrete trajectories closely agree, especially for the smaller learning rate.
(b) Convergence trajectories of $\theta(t)$ with $\theta_{0}=-0.1$ for the nonlocal and discrete models. Mirror behavior of the positive case: convergence to $\theta^{*}\approx-1$ with oscillations for $\beta=0.99$ when $\alpha=0.1$, and smooth, rate-ordered convergence for $\alpha=0.01$.
Figure 10: Convergence trajectories of $\theta(t)$ for the nonlocal and discrete RMSProp models. These two plots show the evolution of the trajectory of $\theta(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ with two initial conditions $\theta_{0}=\pm 0.1$ for RMSProp.
(a) Squared-gradient trajectories of $v(t)$ for the nonlocal and discrete momentum models with positive initialization. Squared-gradient trajectories $v(t)$ with $\theta_{0}=0.1$. For $\alpha=0.1$, nonlocal curves are smoother than the discrete ones; $\beta=0,0.9$ display a transient peak then decay, while $\beta=0.99$ stays small and nearly flat. For $\alpha=0.01$, discrete and nonlocal curves almost overlap: $\beta\in\{0,0.9\}$ show a bell-shaped peak (larger for $\beta=0$) followed by decay, and $\beta=0.99$ remains low throughout.
(b) Squared-gradient trajectories of $v(t)$ for the nonlocal and discrete momentum models with negative initialization. Squared-gradient trajectories $v(t)$ with $\theta_{0}=-0.1$. Same qualitative pattern as for $\theta_{0}=0.1$: nonlocal curves are smoother at $\alpha=0.1$; with $\alpha=0.01$ both models align closely, with the largest peaks for $\beta=0$, smaller for $\beta=0.9$, and consistently low values for $\beta=0.99$.
Figure 11: Squared-gradient trajectories of $v(t)$ for the nonlocal and discrete RMSProp models. These two plots show the evolution of the trajectory of the gradient $v(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ with two initial conditions $\theta_{0}=\pm 0.1$ for RMSProp.
(a) Convergence trajectories of $\theta(t)$ for the nonlocal and discrete Adam models with positive initialization. Convergence of $\theta(t)$ with $\theta_{0}=0.1$. Solid lines: first-order nonlocal continuous Adam; markers: discrete Adam. Two learning rates ($\alpha=0.1,\,0.01$) and parameters $\beta_{1}\in\{0,0.9\}$, $\beta_{2}\in\{0.99,0.999\}$ are considered. With $\alpha=0.1$, the $\beta_{1}=0.9$ configurations exhibit overshooting and lightly damped oscillations, whereas with $\alpha=0.01$ all curves converge smoothly and nearly overlap, approaching $\theta^{*}\approx 1$.
(b) Convergence trajectories of $\theta(t)$ for the nonlocal and discrete Adam models with negative initialization. Convergence of $\theta(t)$ with $\theta_{0}=-0.1$. Same setup as Fig. 12(a). For $\alpha=0.1$, $\beta_{1}=0.9$ produces mild oscillations; for $\alpha=0.01$ the trajectories are smooth and almost identical between the nonlocal and discrete models, converging to $\theta^{*}\!\approx-1$.
Figure 12: Convergence trajectories of $\theta(t)$ for the nonlocal and discrete Adam models. These two plots show the evolution of the trajectory of $\theta(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ with two initial conditions $\theta_{0}=\pm 0.1$ for Adam.
(a) First-moment trajectories of $m(t)$ for the nonlocal and discrete Adam models with positive initialization. First-moment trajectories $m(t)$ with $\theta_{0}=0.1$. At the larger learning rate $\alpha=0.1$, $m(t)$ exhibits damped oscillations around zero—most pronounced for $\beta_{1}=0.9$ and $\beta_{2}=0.999$. With $\alpha=0.01$, all curves show a single hump followed by decay to zero. Nonlocal and discrete results closely match across regimes.
(b) First-moment trajectories of $m(t)$ for the nonlocal and discrete Adam models with negative initialization. First-moment trajectories $m(t)$ with $\theta_{0}=-0.1$. Mirror behavior of the positive case. For $\alpha=0.1$ the oscillations are stronger for $\beta_{1}=0.9$; for $\alpha=0.01$ the profiles are smooth and the nonlocal curves overlap the discrete ones throughout.
Figure 13: Convergence trajectories of $\theta(t)$ for the nonlocal and discrete Adam models. These two plots show the evolution of the trajectory of $\theta(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ with two initial conditions $\theta_{0}=\pm 0.1$ for Adam.
(a) Second-moment trajectories of $v(t)$ for the nonlocal and discrete Adam models with positive initialization. Second-moment trajectories $v(t)$ with $\theta_{0}=0.1$. Peak height and decay depend on the parameters: with $\alpha=0.1$, $\beta_{1}=0.9,\beta_{2}=0.99$ yields the largest transient peaks, whereas $\beta_{2}=0.999$ produces smaller, flatter curves. For $\alpha=0.01$, all settings display the characteristic bell-shaped peak followed by slow decay, and nonlocal and discrete curves nearly coincide.
(b) Second-moment trajectories of $v(t)$ for the nonlocal and discrete Adam models with negative initialization. Second-moment trajectories $v(t)$ with $\theta_{0}=-0.1$. Same qualitative pattern as in Fig. 14(a): larger peaks for $\beta_{2}=0.99$ and much lower values for $\beta_{2}=0.999$. Agreement between nonlocal and discrete models is very close, especially for the smaller learning rate.
Figure 14: Convergence trajectories of $\theta(t)$ for the nonlocal and discrete Adam models. These two plots show the evolution of the trajectory of $\theta(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ with two initial conditions $\theta_{0}=\pm 0.1$ for Adam.
(a) Parameter trajectories of $\theta(t)$ for the nonlocal and discrete AdaGrad models with zero initialization. Theta trajectories (AdaGrad vs. Non-local AdaGrad) for the nonconvex function $\frac{1}{4}(\theta^{2}-1)^{2}$ for the initial value $\theta_{0}=0$.
(b) Gradient trajectories of $\theta(t)$ for the nonlocal and discrete AdaGrad models with zero initialization. Gradient trajectories (AdaGrad vs. Non-local AdaGrad) for the nonconvex function $\frac{1}{4}(\theta^{2}-1)^{2}$ for the initial value $\theta_{0}=0$.
Figure 15: Equilibrium trajectories of $\theta(t)$ and $G(t)$ for the nonlocal and discrete AdaGrad models. These two plots show the evolution of the trajectory of $\theta(t)$ with the gradient $G(t)$ for the nonconvex function $\tfrac{1}{4}(\theta^{2}-1)^{2}$ under AdaGrad. As can be observed, since the system starts directly at the unstable local maximum $\theta_{0}=0$, it remains unchanged. However, even an arbitrarily small perturbation would cause it to converge to one of the local minima. We also present the discrete case, and, as can be seen, the behavior of the continuous system is completely identical.